OpenStudy (anonymous):

Using a directrix of y = -2 and a focus of (2, 6), what quadratic function is created?

pooja195 (pooja195):

@Directrix

pooja195 (pooja195):

Also \(\huge\color{Blue}{WELCOME~TO~OPENSTUDY!!!! }\)

Directrix (directrix):

Do you have equations as options or perhaps just the name of the type curve described? @snugglemuffinman

OpenStudy (anonymous):

yes I equations to choose from they are :

OpenStudy (anonymous):

\[f(x) = -1/8 (x-2) ^ 2-2\]

OpenStudy (anonymous):

f(x) = 1/16 (x-2)^2 -2

OpenStudy (anonymous):

\[f(x) = 1/8 (x-2)^2 -2\]

OpenStudy (anonymous):

\[f(x) = -1/16 (x+2) -2\]

OpenStudy (anonymous):

@Directrix

OpenStudy (jdoe0001):

y = -2 and a focus of (2, 6),

OpenStudy (anonymous):

2?

OpenStudy (jdoe0001):

hmmm 2? is it? how so?

OpenStudy (anonymous):

oh wait no is it 4?

OpenStudy (jdoe0001):

ok.. .how did you get 4?

OpenStudy (anonymous):

counted the distance from -2 to the vertex

OpenStudy (jdoe0001):

do you know what the directrix is? or the focus point? or maybe you need to refresh that section of the chapters first

OpenStudy (jdoe0001):

ohhh... ok.. .so do you know where the vertex is then?

OpenStudy (anonymous):

the distance is 9? from -2 to 6? And the vertex is it at 2?

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

@Destinymasha @nincompoop @pooja195 someone please help..... I will give a medal

OpenStudy (anonymous):

@Abhisar

OpenStudy (jdoe0001):



OpenStudy (jdoe0001):

was a bit lagged, thus

OpenStudy (anonymous):

6 from the 2 marker and 2 from -2 to 2? Sorry im really bad at this